Solve each problem and provide the complete solution.
note: topic is all about Gravitational force

Calculate the gravitational field strength on the surface of Mars. Mars has a radius of 3.43 x 106m and a mass of 6.37x1023kg.
1
Expert's answer
2020-11-30T14:56:40-0500
"g=\frac{GM}{R^2}\\g=\frac{(6.67\cdot10^{-11})(6.37\cdot10^{23})}{(3.43\cdot10^{6})^2}=3.61\frac{m}{s^2}"